find the number of litres that a water tank with length 3 m, breadth 2 m and depth 1.5 m can hold.
Answer by Edwin McCravy(20086)

find the number of litres that a water tank with length 3 m, breadth 2 m and depth 1.5 m can hold.
1m³ = 1000 liters

V = 3m×2m×1.5m = 9m³ = 9×1000 liters = 9000 liters.

Edwin